### v1

@@ 2.4 Categorizing the Identified Articles
After applying the inclusion and exclusion criteria, the remaining publications were carefully reviewed for relevance. Those that did not meet the criteria were excluded from further analysis. The selected papers were then categorized into three distinct groups: methods, datasets, and review papers. Each category was analyzed separately, and the corresponding information was organized into three dedicated tables. Publications relevant to more than one category were included in each applicable table as separate entries. Key information for each paper was systematically recorded in the appropriate table columns, as detailed below:
  • Methods: The name of the proposed method, datasets used for evaluation, a brief description of the architecture, the input modality, the learning strategy, and the evaluation metrics;
  • Datasets: For each dataset, we recorded the following details: name, type of gesture based on temporal characteristics (e.g., static, isolated dynamic, or continuous), data modality, total number of samples, number of subjects, and—when applicable to continuous HGR—the number of gesture instances. For sign language understanding (SLU) datasets, we also noted the language, vocabulary sizes (both sign and text), and total duration (for video-based datasets). Additional attributes include the source of the dataset (e.g., laboratory, television, or web), frame rate, resolution, and the environmental conditions under which the data was recorded.
  • Surveys: For survey papers, a concise description of the content was collected.
(a) The distribution of the selected papers on the top-10 venues, in terms of number of publications.
(b) The number of selected publications grouped by year.
Figure 1: Bibliometric statistics for the selected papers.
Additionally, for each publication, metadata such as title, authors’ names, venue, publication year, citation count, keywords, task, application domain, authors’ motivation, challenges, and future directions were also recorded. This meticulous process resulted in a curated collection of 290 papers. Some bibliometrics are provided in Fig. 1. Notably, the number of review papers is limited, highlighting the research gap that this survey aims to address.
@@ 3 Taxonomy of Hand Gesture Recognition Methods
Figure 2: Taxonomy of hand gesture recognition methods.
This section performs an overview of the landscape of vision-based hand gesture recognition methods, including various aspects, such as input, representation, applications, etc.. More specifically, a set of complementary and diverse criteria, each focusing on key aspects of a VHGR system, is considered, resulting into the classification of related works into a corresponding set of main categories (while each class can be further divided into sub-categories, as will be detailed below). The different criteria that are used to organize the research works in the field are graphically illustrated in Fig. 2 and are detailed as follows:
@@
  • Application Domain: This criterion refers to the various domains in which VHGR (Visual Hand Gesture Recognition) solutions have been applied. The following categories are considered:
@@
    • Sign Language Understanding (SLU): Involves the automatic interpretation of gestures in either a single sign language (monolingual SLU [102, 12]) or across multiple sign languages (multilingual SLU [126]).
@@
    • Human-Computer Interaction (HCI): Encompasses any form of interaction between users and computers. Numerous studies focus on HCI [131, 193], with particular emphasis on AR/VR applications [199, 32, 5].
@@
    • Human-Robot Interaction (HRI): Includes gesture-based control of mobile robots such as UGVs [17, 152] and UAVs [211, 194], industrial robotic arms [96], assistive robots [29], and surgical robots [139].
@@
    • Human-Machine Interaction (HMI): Focuses on interfaces between humans and machines, especially in automotive contexts. VHGR enables recognition of traffic gestures (e.g., from police officers) [45, 59, 185] and supports the development of more intuitive in-vehicle user interfaces [124, 138, 53, 115, 52].
@@
  • Task: Complementing the application domains, this criterion focuses on the nature of the gestures themselves, specifically their temporal characteristics. Gestures are broadly classified as either static or dynamic, depending on whether they involve a fixed hand posture or a sequence of movements over time. Dynamic gestures typically consist of three temporal phases: preparation, core (or nucleus), and retraction [124], which reflect the gesture’s evolution from initiation to completion. Illustrative examples are provided in Fig. 3. The recognition of dynamic gestures can be further divided into isolated and continuous, considering only one gesture per data sequence and identifying gestures in a continuous manner, respectively.
    Figure 3: Gesture recognition based on task: a) static [33] and b) dynamic [27] VHGR.
@@
  • Input Modality: In continuation of the previous criteria, which focused on application contexts and gesture types, this criterion examines the nature and variety of visual input data used for VHGR. The choice of input modality directly influences the representation architecture and the deep learning models employed. While RGB frames remain the most commonly used modality, other forms of visual data have also been explored—such as depth maps, infrared (IR) images, skeleton joints (and derived features like bones and motion [38]), pose heatmaps [27, 214], optical flow, and motion history [119]. Representative examples are shown in Fig. 4.
    Based on the type of input, methods are broadly categorized as:
    • Pose-based (or skeleton-based): These approaches exclusively process skeleton data.
    • Appearance-based: These rely on visual cues from RGB, depth, IR, or similar modalities.
@@
    (a) From left to right: RGB, Optical Flow, Depth and IR [124]
    (b) Skeleton joints [102]
    (c) Pose heatmap [27]
    (d) Motion history [119]
@@
    In terms of the number of input modalities, methods can be classified as:
    • Uni-modal: Utilize a single type of input data.
    • Multi-modal: Integrate two or more modalities. These methods can be further categorized based on their multi-modality fusion strategy [177, 60] (see Fig. 5):
        Early (data-level) fusion: Combines modalities at the input level before processing [49, 180].
        Late (score-level) fusion: Fuses final prediction outputs from independent uni-modal models [119, 167, 51].
        Intermediate (feature-level) fusion: Merges features extracted from separate modalities at a single processing level  [172].
        Multi-level fusion: Enables interaction across modalities at multiple stages of processing [214, 60].
    Figure 5: Multi-modality fusion techniques: a) Early (data-level) fusion, b) Late (score-level) fusion, c) Intermediate (feature-level) fusion, and d) Multi-level fusion. Illustration from [60]
  • Type of Information Conveyed: Building upon the previous criterion, which focused on the nature of the input, this dimension considers the meaning or intent behind the gestures. Specifically, it examines the kind of information gestures are meant to transmit [16], resulting in the following categories (illustrated in Fig. 6):
    • Semaphoric gestures convey specific commands or instructions and are commonly used in gesture-based interfaces [124, 131, 17].
    • Language gestures function similarly to semaphoric ones but are exclusively used in the context of sign languages [12, 158].
    • Deictic (pointing) gestures are used to indicate objects, directions, or spatial targets [127].
    • Manipulative gestures aim to interact with or mimic the manipulation of physical objects [50].
    (a) Semaphoric [86]
    (b) Language [12]
    (c) Pointing [127]
    (d) Manipulative [50]
@@
  • Camera Point of View: Complementing the previous criteria on gesture content and input modalities, this dimension focuses on the spatial relationship between the camera and the performer [199]. The camera’s perspective influences both the visual characteristics of the gesture and the recognition model design. Three primary viewpoints are considered (see Fig. 7):
@@
    (a) First-person view (egocentric) [199]
    (b) Second-person view [149]
    (c) Third-person view [158]
@@
  • Architecture: Building on the previous criteria, which focused on the nature of input and camera perspective, this criterion concerns the architectural design of the gesture recognition system, regarding the modeling stages performed by the backbone. Specifically, it refers to how spatial and temporal information is processed and integrated to model gesture dynamics. Three main categories are identified:
@@
  • Learning Strategy: This criterion explores the different training methodologies adopted by researchers to develop VHGR models, often influenced by data availability, task complexity, and desired generalization. The main strategies identified include:
@@
    • Transfer Learning, where models are pretrained on large-scale datasets (e.g., Jester [117]) and subsequently fine-tuned for specific downstream VHGR tasks [112].
@@
    • Weakly Supervised Learning, often used in Continuous Hand Gesture Recognition (CHGR), where temporal annotations are coarse or unaligned, lacking exact gesture boundaries [188].
@@
    • Knowledge Distillation, in which one part of the network (often a teacher model) transfers knowledge to another (student model) to enhance feature extraction and efficiency [122, 197, 27].
@@
    • Contrastive Learning, used either in a CLIP-based framework to support zero-shot/few-shot recognition by leveraging textual descriptions [22, 146, 23], or to improve the discriminative capacity of learned representations [203, 204].
@@
    • Multi-task Learning, where a single model is optimized to perform multiple related tasks simultaneously, encouraging knowledge sharing and improved generalization [102].
@@
    • Reinforcement Learning, applied to optimize decision-making processes, such as keyframe selection, by rewarding effective representations [46].
@@
    • Few/Zero-shot Learning, which enable the model to generalize to new gestures or languages with limited or no annotated examples, often leveraging semantic or textual guidance [22, 146, 23].
@@
    • Adversarial Learning, used as a data augmentation strategy [13] to improve robustness, as demonstrated in adversarial video synthesis networks [128].
@@
    • Masked Reconstruction Pretraining, inspired by BERT-style architectures, in which models are pretrained to reconstruct masked skeleton sequences, injecting structural prior knowledge (e.g., hand topology) [202, 71, 68].
This comprehensive taxonomy of vision-based hand gesture recognition (VHGR) methods, has been grounded in a diverse set of classification criteria that span from application domains to architectural and training strategies. Among these, the task criterion, distinguishing between static and dynamic gestures based on their temporal properties, emerges as particularly pivotal. It fundamentally shapes not only the selection of appropriate models and input modalities but also the interpretation of gestures in real-world scenarios. Due to its central role in determining the complexity and design of recognition systems, the task criterion serves as a natural point of departure for further analysis. The next three sections delve deeper into this dimension by systematically examining the core types of VHGR tasks—Static Visual Hand Gesture Recognition, Isolated Dynamic VHGR, and Continuous VHGR—each with its own challenges, methodologies, and applications.
@@ 4 Methods for Static VHGR
Static VHGR aims to identify gestures relying only on spatial data. As this approach neglects temporal information, is only suitable for recognizing static hand gestures. Proposed methods for static VHGR can be further categorized into single-stage and multi-stage approaches, based on whether the architecture consists of a single model or a multi-model processing pipeline, respectively. Indicative works of each category are briefly analyzed into the two following subsections. The third subsection concludes with the main features of state-of-the-art static VHGR methods, and discusses the open challenges that concern the community.
4.1 Single-stage Approaches
@@
Single-stage approaches rely on a single model, which varies depending on the task. For gesture classification, fine-tuned variants of well-known 2D-CNN architectures are employed, such as VGG and DenseNet, or custom lightweight architectures are trained from scratch. When classification is combined with localization of gestures, variants of YOLO are adopted. As far as it concerns the application, static VHGR methods are used in Sign Language Understanding (SLU) and Human Computer Interaction (HCI). However, a few more general-purpose models have also been proposed:

### v2

@@ 2.4 Categorizing the Identified Articles
After applying the inclusion and exclusion criteria, the remaining publications were carefully reviewed for relevance. Those that did not meet the criteria were excluded from further analysis. The selected papers were then categorized into three distinct groups: methods, datasets, and review papers. Each category was analyzed separately, and the corresponding information was organized into three dedicated tables. Publications relevant to more than one category were included in each table as separate entries. This procedure resulted in 185 articles presenting methods, 72 presenting datasets, and 21 survey articles. Key information for each paper was systematically recorded in three long tables which contain the appropriate table columns, as detailed below:
  • 1.
    Methods: For each article, we recorded the name of the method it proposes, the datasets used for evaluation, the evaluation metrics, a brief description of the architecture, the input modality, and the learning strategy.
  • 2.
    Datasets: For each dataset, we recorded its name, the type of gesture based on temporal characteristics (e.g., static, isolated dynamic, or continuous), the data modality, the total number of samples, the number of subjects, and—when applicable to CHGR—the number of gesture instances. For Sign Language Understanding (SLU) datasets, we also noted the language, vocabulary sizes (both sign and text), and total duration (for video-based datasets). Additional attributes include the source of the dataset (e.g., laboratory, television, or web), frame rate, resolution, and the environmental conditions under which the data was recorded.
  • 3.
    Surveys: For survey papers, we provide a concise description of the content. Furthermore, the proposed taxonomy, the time range covered, and the scope of the survey were recorded. Although we examine reviews considering both sensor-based and vision-based HGR techniques, we focus only on content related to VHGR.
For each publication, several metadata such as title, authors’ names, venue, publication year, citation count, keywords, task, application domain, authors’ motivation, challenges, and future directions were recorded. The tables are not presented directly in this article, due to their size, but are available upon request. The most recent surveys have already been presented in Table 1 and discussed in the Introduction section, the methods and datasets are presented in an organised manner in the following Sections, and some bibliometrics are provided in Fig. 1.
Papers Range
Methods 2021-2025
Datasets 2010-2025
Reviews 2021-2025
Figure 1: Bibliometric statistics. Figure a) shows the distribution of the selected papers on the top-10 venues, in terms of number of publications. Almost 50 of the papers considered in the current survey have been published in IEEE Access, covering the whole spectrum of domains and tasks. Figure b) illustrates the number of publications & pre-prints per year. As the current survey focuses on recent advancements, most of the selected papers have been published since 2021. Methods and reviews published before 2021 were ignored. The number of methods, datasets and reviews (which may be overlapped) are shown in figure c). The publication range for each category is reported in table d).
@@ 3 Taxonomy of Hand Gesture Recognition Methods
Figure 2: Taxonomy of hand gesture recognition methods. A diverse set of criteria has been proposed by many works. This survey adds a few more, including ”Architecture”, considering it essential for a method-oriented presentation.
To better understand the diversity and evolution of vision-based hand gesture recognition (VHGR) methods, this section introduces a structured taxonomy of existing approaches. It classifies related works according to complementary criteria that capture key aspects of VHGR systems—such as input type, data representation, and application domain—each further divided into subcategories. The overall taxonomy is illustrated in Fig. 2 and described in detail below.
3.1 Application Domain
@@
This criterion refers to the various domains in which VHGR solutions have been applied. The following categories are considered:
  • 1.
@@
    Sign Language Understanding (SLU): Involves the automatic interpretation of gestures in either a single sign language (monolingual SLU (UNISIGNli2025; BOBSL-BBC)) or across multiple sign languages (multilingual SLU (SIGN-muller-etal-2023-findings)).
  • 2.
@@
    Human-Computer Interaction (HCI): Encompasses any form of interaction between users and computers. Numerous studies focus on HCI (HAGRIDv2_ARXIV; ZJUGesture), with particular emphasis on AR/VR applications (EgoGesture_DATASET; OO-dMVMT; BOAT-MI_VR).
  • 3.
@@
    Human-Robot Interaction (HRI): Includes gesture-based control of mobile robots such as UGVs (ULTRA-RANGE-UGV-Bamani; MOBILE_ROBOT_DRCAM) and UAVs (LONG_RANGE_HGR_SSD; MD-UHGRD_SA-YOLO), industrial robotic arms (ROBOTIC_ARM), assistive robots (DIALOG_POINTING), and surgical robots (TELEOPERATED_SURGICAL_ROBOT).
  • 4.
@@
    Human-Machine Interaction (HMI): Focuses on interfaces between humans and machines, especially in automotive contexts. VHGR enables recognition of traffic gestures (e.g., from police officers) (FA-STGCN; MG-GCT; CTCX_WANG2022123) and supports the development of more intuitive in-vehicle user interfaces (NVIDIA_GESTURE_R3DCNN_Molchanov; EgoFormer_METHOD_EgoDriving_DATASET; ConvMixFormer; Briareo_dataset; MVTN_ARXIV).
3.2 Task
@@
Complementing the application domains, this criterion focuses on the nature of the gestures themselves, specifically their temporal characteristics. Gestures are broadly classified as either static or dynamic, depending on whether they involve a fixed hand posture or a sequence of movements over time. Dynamic gestures typically consist of three temporal phases: preparation, core (or nucleus), and retraction (NVIDIA_GESTURE_R3DCNN_Molchanov), which reflect the gesture’s evolution from initiation to completion. The recognition of dynamic gestures can be further divided into isolated and continuous, considering only one gesture per data sequence and identifying gestures in a continuous manner, respectively.
3.3 Input
@@
In continuation of the previous criteria, which focused on application contexts and gesture types, this criterion examines the nature and variety of visual input data used for VHGR. The choice of input modality directly influences the representation architecture and the deep learning models employed. While RGB frames remain the most commonly used modality, other forms of visual data have also been explored—such as depth maps (including RGB-like representations, such as HHA (SAM_SLR)), infrared (IR) images, skeleton joints (and derived features like bones and motion (TMS-NET)), pose heatmaps (NEURIPS2022_6cd3ac24_HEATMAPS; NLA_SLR), optical flow and motion history (MOTION_HISTORY_Mercanoglu).
In terms of the number of input modalities, methods are classified as uni-modal, when they use a single type of input data, or multi-modal, when they integrate two or more modalities. In the case of multiple modalities in the input, a multi-modality fusion strategy is chosen (MMTM; CTFB_Hampiholi). This could be:
  • 1.
    Early (data-level) fusion: Combines modalities at the input level before processing (DHGR_3D_POSE_HRI; 3STREAM_DISTILLATION).
  • 2.
    Late (score-level) fusion: Fuses final prediction outputs from independent uni-modal models (MOTION_HISTORY_Mercanoglu; JOINTS_MOTION_VECTORS_Improved_ST-GCN; GestFormer).
  • 3.
    Intermediate (feature-level) fusion: Merges features extracted from separate modalities at a single processing level  (tang2021graph).
  • 4.
    Multi-level fusion: Enables interaction across modalities at multiple stages of processing (NLA_SLR; CTFB_Hampiholi).
Finally, based on the type of input, the methods are broadly categorized into two groups: pose-based (or skeleton-based) approaches, which exclusively process skeleton data, and appearance-based approaches, which rely on visual cues from RGB, depth, infrared, or similar modalities.
@@
(a) From left to right: RGB, Optical Flow, Depth and IR (NVIDIA_GESTURE_R3DCNN_Molchanov)
(b) Skeleton joints (UNISIGNli2025)
(c) Pose heatmap (NEURIPS2022_6cd3ac24_HEATMAPS)
(d) Motion history (MOTION_HISTORY_Mercanoglu)
@@
3.4 Type of Information Conveyed
Building upon the previous criterion, which focused on the nature of the input, this dimension considers the meaning or intent behind the gestures. Specifically, it examines the kind of information gestures are meant to transmit (TAXONOMY_OF_HAND_GESTURES_2000_ASL_DATASET), resulting in the following categories (illustrated in Fig. 4):
  • 1.
    Semaphoric gestures convey specific commands or instructions and are commonly used in gesture-based interfaces (NVIDIA_GESTURE_R3DCNN_Molchanov; HAGRIDv2_ARXIV; ULTRA-RANGE-UGV-Bamani).
  • 2.
    Language gestures function similarly to semaphoric ones but are exclusively used in the context of sign languages (BOBSL-BBC; Auslan-Daily).
  • 3.
    Deictic (pointing) gestures are used to indicate objects, directions, or spatial targets (DeePoint).
  • 4.
    Manipulative gestures aim to interact with or mimic the manipulation of physical objects (FPHA).
(a) Semaphoric (WACV_kapitanov2024hagrid)
(b) Language (BOBSL-BBC)
(c) Pointing (DeePoint)
(d) Manipulative (FPHA)
@@ 3.4 Type of Information Conveyed
3.5 Camera Point of View
Complementing the previous criteria on gesture content and input modalities, this dimension focuses on the spatial relationship between the camera and the performer (EgoGesture_DATASET). The camera’s perspective influences both the visual characteristics of the gesture and the recognition model design. Three primary viewpoints are considered (see Fig. 5):
  • 1.
@@ 3.5 Camera Point of View
(a) First-person view (egocentric) (EgoGesture_DATASET)
(b) Second-person view (Lsa64argentiniansignlanguage)
(c) Third-person view (Auslan-Daily)
@@
3.6 Architecture
Building on the previous criteria, which focused on the nature of input and camera perspective, this criterion concerns the architectural design of the gesture recognition system, regarding the modeling stages performed by the backbone. Specifically, it refers to how spatial and temporal information is processed and integrated to model gesture dynamics. Three main categories are identified:
  • 1.
@@ 3.7 Learning Strategy
This criterion explores the different training methodologies adopted by researchers to develop VHGR models, often influenced by data availability, task complexity, and desired generalization. The main strategies identified include:
  • 1.
@@
    Transfer Learning, where models are pretrained on large-scale datasets (e.g., Jester (Jester-Dataset)) and subsequently fine-tuned for specific downstream VHGR tasks (pretrained_on_jester_3dDenseNet_LSTM).
  • 2.
@@
    Weakly Supervised Learning, often used in Continuous Hand Gesture Recognition (CHGR), where temporal annotations are coarse or unaligned, lacking exact gesture boundaries (Fangyun_Wei_ICCV).
  • 3.
@@
    Knowledge Distillation, in which one part of the network (often a teacher model) transfers knowledge to another (student model) to enhance feature extraction and efficiency (VAC-CSLR; C2ST; NEURIPS2022_6cd3ac24_HEATMAPS).
  • 4.
@@
    Contrastive Learning, used either in a CLIP-based framework to support zero-shot/few-shot recognition by leveraging textual descriptions (ZERO-SHOT-SLR; ZS-GR; FEW-SHOT-BILGE2024110374), or to improve the discriminative capacity of learned representations (MASA; Weichao-Zhao).
  • 5.
@@
    Multi-task Learning, where a single model is optimized to perform multiple related tasks simultaneously, encouraging knowledge sharing and improved generalization (UNISIGNli2025).
  • 6.
@@
    Reinforcement Learning, applied to optimize decision-making processes, such as keyframe selection, by rewarding effective representations (KEYFRAMES_DRL_SKELETON).
  • 7.
@@
    Few/Zero-shot Learning, which enable the model to generalize to new gestures or languages with limited or no annotated examples, often leveraging semantic or textual guidance (ZERO-SHOT-SLR; ZS-GR; FEW-SHOT-BILGE2024110374).
  • 8.
@@
    Adversarial Learning, used as a data augmentation strategy (alimisis2025advances) to improve robustness, as demonstrated in adversarial video synthesis networks (ADVERSARIAL_DATA_AUGM).
  • 9.
@@
    Masked Reconstruction Pretraining, inspired by BERT-style architectures, in which models are pretrained to reconstruct masked skeleton sequences, injecting structural prior knowledge (e.g., hand topology) (BEST; SignBERT+; SignBERT_ICCV).
3.8 Processing
Existing methods for isolated and continuous dynamic HGR can be further categorized based on whether the models have full access to the entire gesture(s) sequence or not. According to (NVIDIA_GESTURE_R3DCNN_Molchanov), two types of processing strategies apply:
  • 1.
    Online (real-time) processing: Online HGR models perform gesture recognition in a stream of input frames. In other words, these models take as input the gesture sequence recorded until a specific time. Thus, online gesture recognition is suitable for real-time applications, as the gesture is recognized before the retraction part, reducing the latency. Sliding windows and gesture detector/classifier pipelines are often employed (IPN-HAND; survey-emporio2025continuous).
  • 2.
    Offline: Offline processing means that the model treats the entire gesture sequence. This approach is limited to applications that do not require real-time recognition. Most of the existing methods for VHGR fall into this category, such as those using BiLSTM for temporal modeling (VAC-CSLR).
The presented taxonomy of VHGR methods highlights the diversity of approaches developed to address different aspects of gesture understanding—from isolated dynamic gestures to sign language recognition and multimodal human–computer interaction. It provides a comprehensive foundation for analyzing current trends and identifying research gaps. Building on this overview, the following three sections delve deeper into specific VHGR categories, examining in detail the latest methods for Isolated Dynamic, Continuous, and Multimodal gesture recognition, respectively.
@@ 4 Methods for Static VHGR
Static VHGR (SHGR) aims to identify gestures relying only on spatial data. As this approach neglects temporal information, is only suitable for recognizing static hand gestures. For gesture classification, fine-tuned variants of well-known 2DCNN architectures are employed, such as VGG and DenseNet, or custom lightweight architectures are trained from scratch. When classification is combined with localization of gestures, variants of YOLO are adopted. Hand segmentation is often applied to eliminate the background, reducing the interference of gesture-irrelevant information. Hand detection is also utilized in some cases to normalize the spatial dimensions, especially for long-range recognition. As far as it concerns the application, static VHGR methods are used in SLU, HCI and HRI. However, a few more general-purpose models have also been proposed.
4.1 Sign Language Understanding (SLU)